24 inch wire, cut into 2 piece, twice the length of shorter piece is equal to the length of the longer piece, what is the length of the longer piece?
Answer by checkley77(12844)

X+Y=24 OR X=24-Y
2X=Y
X=24-2X
X+2X=24
3X=24
X=24/3
X=8 ANS. FOR THE SHORTER LENGTH OF WIRE.
2*8=16 ANS. FOR THE LONGER PIECE OF WIRE.
PROOF:
8+16=24
24=24